countryboy3642:

Solve for x. x3=−1000 Enter your answer in the box. x =

Pixel:

-333.3333333333333333333333333

harliii:

the answer is -10

Pixel:

-10 x 3 = -30 -_-

countryboy3642:

it is r u sure

harliii:

10 to the third power

Pixel:

no there is no x^3

countryboy3642:

is it -30 or -10

Pixel:

Country is the 3 above the x?

harliii:

are u squared the 3 are what

countryboy3642:

x3=-1000

Pixel:

\[x ^{3}\]

Pixel:

does it look like that?

mplatt:

you sure its not 3x=1000 instead of x3=1000?

mplatt:

if so Let's solve your equation step-by-step. 3x=1000 Step 1: Divide both sides by 3. 3x/3=1000/3 x= 1000/3 Answer: x= 1000 3